y'y'''-(y'')^2=(y'')^3


1
Expert's answer
2021-05-07T10:12:23-0400


Solution :-


y'y'''-(y'')^2=(y'')^3


= "({dy\over dx})({d^3y\over dx^3})-({d^2y\over dx^2})^2=({d^2y\over dx^2})^3"



="({dy\over dx})({d^3y\over dx^3})-({d^2y\over dx^2})^2-({d^2y\over dx^2})^3 = 0" ........(1)

manuplating the (1) equation



"f\left( x \right) \cdot {\large\frac{d}{{dx}}\normalsize} \left[ {\ln f\left( x \right)} \right]" = "\frac {dy}{dx}" .....(2)



= "\intop\intop\intop F(y)dx"


x = log(y)

using the properties and (1) , (2)

y = ex


="\intop\intop C_1e^{xc_1}.e^{c_1}" dx.dx



="\intop e^{xc_1}.e^{c_1c_2}" dx


answer = "\boxed{y(x)={e^{xc_1}e^{c_1c_2}\over c_1} + C_3}"